A man made a deposit of P1060. iF HIS DEPOSIT CONSISTED OF 42 BILLS, SOME 10PESO BILLS AND THE REST 50 PESO BILLS, HOW MANY OF BILLS OF EACH KIND DID HE HAVE?
Answer by checkley71(8403)

T+F=42 OR T=42-F
10T+50F=1060
10(42-F)+50F=1060
420-10F+50F=1060
40F=1060-420
40F=640
F=640/40
F=16 NUMBER OF 50 PESO BILLS.
42-16=26 NUMBER OF 10 PESO BILLS.
PROOF
10*26+50*16=1060
260+ 800=1060
1060=1060